# On robust overfitting: adversarial training induced distribution matters

### 28 Nov 2023 v1

@@ Abstract
Adversarial training may be regarded as standard training with a modified loss function. But its generalization error appears much larger than standard training under standard loss. This phenomenon, known as robust overfitting, has attracted significant research attention and remains largely as a mystery. In this paper, we first show empirically that robust overfitting correlates with the increasing generalization difficulty of the perturbation-induced distributions along the trajectory of adversarial training (specifically PGD-based adversarial training). We then provide a novel upper bound for generalization error with respect to the perturbation-induced distributions, in which a notion of the perturbation operator, referred to “local dispersion”, plays an important role.
@@ 1 Introduction
Despite their outstanding performance, deep neural networks (DNNs) are known to be vulnerable to adversarial attacks where a carefully designed perturbation may cause the network to make a wrong prediction [1, 2]. Many methods have been proposed to improve the robustness of DNNs against adversarial perturbations [3, 4, 5], among which Projected Gradient Descend based Adversarial Training (PGD-AT) [3] is arguably the most effective [6, 7]. A recent work in [8] however revealed a surprising phenomenon in PGD-AT: after training, even though the robust error (i.e., error probability in the predicted label for adversarially perturbed instances) is nearly zero on the training set, it may remain very high on the testing set. For example, on the testing set of CIFAR10, the robust error of PGD-AT trained model can be as large as 44.19%. This significantly contrasts the standard training: on CIFAR10, when the standard error (i.e., the error probability in the predicted label for non-perturbed instances) is nearly zero on the training set, its value on the testing set is only about 4%. This unexpected phenomenon arising in PGD-AT is often referred to as robust overfitting.
@@
(a) WRN-34
(b) PRN-18
Figure 1: PGD-AT and the corresponding IDE results of the the Wide ResNet34 (WRN-34) model [18] and pre-activation ResNet18 (PRN-18) model [19] on the CIFAR-10 dataset. A significant increase in the IDE testing error is observed on both figures with the appearance of robust overfitting, suggesting a correlation between the generalization difficulty on $\tilde{\mathcal{D}}_{t}$ and robust overfitting.
@@ 3 Adversarial training and induced distributions
We consider a classification setting with input space ${\mathcal{X}}\subseteq\mathbb{R}^{d}$ and label space ${\mathcal{Y}}:=\{1,2,\cdots,K\}$. We will use $\mathcal{D}$ to denote the data distribution on $\mathcal{X}\times\mathcal{Y}$. Let $\Theta\subseteq\mathbb{R}^{n}$ be the model parameter space. For every $\theta\in\Theta$, let $l_{\theta}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ denote a loss function, where $l_{\theta}(x,y)$ is the loss (e.g, the cross-entropy loss) of $(x,y)$ for the model with parameter $\theta$. The adversarial population error $R_{\mathcal{D}}^{\rm adv}(\theta)$ is then
@@
where we have chosen $\mathbb{B}(x,\epsilon):=\{t\in\mathbb{R}^{d}:\|t-x\|_{\infty}\leq\epsilon\}$ as the $\infty$-norm ball around $x$ with the radius $\epsilon$.
In adversarial training, the ultimate objective is to find a model parameter $\theta$ that minimizes $R_{\mathcal{D}}^{\rm adv}(\theta)$. Having no access to $\mathcal{D}$, in practice, a natural choice is to minimize an empirical version of $R_{\mathcal{D}}^{\rm adv}(\theta)$, namely, on a training set $S:=\{(x_{i},y_{i})\}_{i=1}^{m}$ drawn i.i.d from $\mathcal{D}$, solve
$\min\limits_{\theta\in\Theta}R_{S}^{\rm adv}(\theta),{\quad\rm where\quad}R_{S}^{\rm adv}(\theta):=\frac{1}{m}\sum\limits_{i=1}^{m}\max\limits_{v_{i}\in\mathbb{B}(x_{i},\epsilon)}l_{\theta}(v_{i},y_{i})$ (2)
The most popular adversarial training technique of solving this problem for neural networks is iterating between solving the inner maximization via $k$-step projected gradient descend (PGD) and updating $\theta$ through stochastic gradient descent. We now give a concise explanation of this procedure, referred to as PGD-AT, restating the procedure in [3].
@@
Here $\Pi_{\mathbb{B}(x,\epsilon)}:\mathbb{R}^{d}\to\mathbb{B}(x,\epsilon)$ denotes the operation of projecting onto the set $\mathbb{B}(x,\epsilon)$ and $\lambda\in\mathbb{R}_{+}$ is a hyperparameter. The sgn($\cdot$) function takes the sign of a vector on an element-wise basis. The $k$-step PGD mapping ${\mathcal{Q}}_{x,y,\theta}:\mathbb{R}^{d}\to\mathbb{B}(x,\epsilon)$ is then
@@
For later use, we denote the perturbation family $\{{\cal Q}_{x,y,\theta}:(x,y)\in{\cal X}\times{\cal Y}\}$ by ${\cal Q}_{\theta}$.
Iterations of PGD-AT In PGD-AT, the process of generating a perturbed example $(v,y)$ from an example $(x,y)$ w.r.t a model parameter $\theta$ can be described as
$v$ $={\mathcal{Q}}_{x,y,\theta}(x+\rho)$ (5)
where $\rho$ is drawn from $\mathcal{U}([-\epsilon,+\epsilon]^{d})$, the uniform distribution over the $d$-dimensional cubic $[-\epsilon,+\epsilon]^{d}$. At iteration $t$, where the model parameter is $\theta_{t}$, the solution of the inner maximization $\max_{v\in{\mathbb{B}}(x_{i},\epsilon)}l_{\theta_{t}}(x_{i},y_{i})$ is taken as $l_{\theta_{t}}(v_{i},y_{i})$, and the model parameter is updated by
$\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta_{t}}\left[\frac{1}{m}\sum\limits_{i=1}^{m}l_{\theta_{t}}\left(v_{i},y_{i}\right)\right]$ (6)
It is worth noting that the stochastic mapping (5) that perturbs $(x,y)$ to $(v,y)$ depends on $\theta$. Thus the distribution of the random variable pair $(v,y)=({\mathcal{Q}}_{x,y,\theta_{t}}(x+\rho),y)$ at iteration $t$ depends on $\theta_{t}$. We will denote this distribution by $\tilde{\cal D}_{\theta_{t}}$, or simply $\tilde{\cal D}_{t}$, and refer to it as the (adversarial training) induced distribution at iteration $t$. Then at iteration $t$, we may regard the perturbed examples $\{(v_{i},y_{i})\}$ as an i.i.d. sample from $\tilde{\cal D}_{t}$.
Note that the distribution $\tilde{\mathcal{D}}_{t}$ evolves with $\theta_{t}$ during training and in turn affects the update of $\theta_{t}$. Since perturbation at every iteration is restricted to a small neighborhood of $x$, $\tilde{\mathcal{D}}_{t}$ is “not far” from the original data distribution $\mathcal{D}$. However, in our following experiments, we show that the induced distributions $\tilde{\mathcal{D}}_{t}$ can be much harder to generalize than the original distribution $\mathcal{D}$, particularly when robust overfitting occurs.
@@ 4 Training on the induced distributions
The following experiments are conducted. First PGD-AT is performed on a training set $S$. Along this process, for a prescribed set of training iterations (or “checkpoints”) $I:=\{k_{1},k_{2},\ldots,k_{N}\}$, the model parameter $\theta_{t}$ at each checkpoint $t\in I$ is saved. Then at each checkpoint $t$, each example in the training set $S$ is perturbed according to (5) with $\theta=\theta_{t}$, giving rise to the perturbed training set $\tilde{S}_{t}$. The testing set $T$ is also similarly perturbed, giving rise to perturbed testing set $\tilde{T}_{t}$. The model is then retrained fully on $\tilde{S}_{t}$, using standard training (i.e, without perturbation) with random initialization. The resulting model is tested on $\tilde{T}_{t}$. Note that in this setting, both $\tilde{S}_{t}$ and $\tilde{T}_{t}$ are i.i.d. samples from $\tilde{\cal D}_{t}$. We call these experiments “induced distribution experiment” (IDE) for the ease of reference.
@@
(a) CIFAR-100
(b) Reduced Imagenet
Figure 2: PGD-AT and the corresponding IDE results on CIFAR-100 and the reduced ImageNet. The behaviour of the red curves matches that of the yellow curve, as we observe a substantial rise in IDE testing errors concurrent with the emergence of robust overfitting. This demonstrate a compelling correlation between these two quantities.
(a) MNIST
(b) CIFAR-10 (weight decay)
@@
Figure 3: (a)PGD-AT and the corresponding IDE results on MNIST. The experiment presents a scenario that a good robust generalization is achieved. The absence of robust overfitting coincides with the consistently low IDE testing error. (b) exhibits the outcomes of additional experiments conducted on CIFAR-10. In the experiments, we perform PGD-AT with various weight decay rates and conduct IDEs for each of the PGD-AT variant. The blue curves are reproduced from Figure1 (a), serving as a reference for a clear comparison. The results further solidify the correlation between the robust overfitting and the IDE testing error.
@@
For each dataset, the experiments are repeated five times with different random seeds. The experimental results on CIFAR-10 and CIFAR-100 are shown in Figure 1 and 2 (a), where the green and yellow curves illustrate a phenomenon known as robust overfitting [8]: after a certain point in PGD-AT, the robust generalization gap (yellow curves), that is $|R_{\mathcal{D}}^{\rm rob}(\theta_{t})-R_{S}^{\rm rob}(\theta_{t})|$, steadily increases while the adversarial training error (green curves) constantly decreases. The red curves in the figures depict the testing errors of IDEs w.r.t various checkpoints in PGD-AT. Notably, a significant rise in the average IDE testing error is observed between the 80th and 120th checkpoints, increasing from 3.6% to 23.89% for CIFAR-10 with WRN and from 19% to 48.99% for CIFAR-100. Furthermore, this shift coincides with the onset of robust overfitting, where a significant rise in robust generalization gap is also observed. These results indicate that as $\theta_{t}$ evolves along PGD-AT, the model trained on the samples of $\tilde{\mathcal{D}}_{t}$ becomes harder to generalize. More importantly, this difficulty in generalization appears to be closely linked to the robust overfitting phenomenon. This potential connection is further demonstrated by the experimental results on the reduced ImageNet (see Figure2 (b)), where robust overfitting emerges at an earlier training stage and simultaneously a rise in the IDE testing error occurs. This increment in the IDE testing error is also substantial, with an averaged error of 21.65% at the 20th checkpoint elevating to 38.52% at the 60th checkpoint.
@@
Our experiments on MNIST (see Figure 3 (a)) exhibits a scenario where a good robust generalization is achieved. Interestingly, a small IDE testing error is maintained throughout the evolution of $\tilde{\mathcal{D}}_{t}$ with the absence of robust overfitting. Figure 3 (b) shows results from additional experiments on CIFAR-10. In these experiments, we perform PGD-AT with different level of weight decay to control the level of robust overfitting. Subsequently, IDEs are conducted for each such variant of PGD-AT. In Figure 3 (b), each distinct color corresponds to a different weight decay factor utilized in PGD-AT. Within each color category, the dashed curves and the corresponding solid lines represent, respectively, the robust generalization gaps and the IDE results associated with that specific PGD-AT variant. As anticipated, increasing the weight decay factor results in a notable reduction in the robust generalization gap, while conversely, decreasing the weight decay factor leads to the opposite effect. This is shown by the downward shift in the dashed curves across the three color categories. Additionally, a clear synchronization can be observed between each pair of dashed and solid curves, with lower dashed curves consistently corresponding to lower solid curves in the same color category.
@@
At this end, we have established that the increasing difficulty of generalization inherent in the induced distribution plays an important role in robust overfitting. These observations highlight the impact of the dynamics of PGD-AT on robust overfitting, beyond the static quantities, such as loss landscape.
@@ 5 Generalization properties of the induced distributions
There are two views of adversarial examples, “in-distribution view” (e.g., in [44, 45, 46]) and ”out-of-distribution view” (e.g., in [1, 47, 48]). In the out-of-distribution view, adversarial examples are considered as living outside the data manifold. In the in-distribution view, adversarial examples are considered as located within the support of the true data distribution but having low probability (density). This paper takes the in-distribution view and make the following assumption on the data distribution ${\cal D}$, which we may re-express as the pair $({\cal D}_{\cal X},h^{*})$, where ${\cal D}_{\cal X}$ is the marginal distribution of ${\cal D}$ on the input space ${\cal X}$ and $h^{*}$ is the ground-truth classifier: There is a distribution ${\cal D}_{\cal X}^{*}$ on ${\cal X}$ and some small $\epsilon>0$, such that for every $x$ in the support of ${\cal D}^{*}_{\cal X}$, $h^{*}(x)=h^{*}(x+\rho)$ for all $\|\rho\|_{\infty}\leq\epsilon$, and $D_{\cal X}$ is the convolution of ${\cal D}^{*}_{\cal X}$ with the uniform distribution on ${\mathbb{B}}(0,\epsilon)$. For the simplicity of notation, we write the $(x,h^{*}(x))$ as $(x,y)$, and when $x$ is drawn from ${\cal D}^{*}_{\cal X}$, we denote the joint distribution of $(x,y)$ by ${\cal D}^{*}$.
@@
We now consider a standard classification problem with $\tilde{\mathcal{D}}_{\theta}$ as the underlying data distribution over $\mathcal{X}\times\mathcal{Y}$. Let ${\cal H}$ be a hypothesis class for this learning problem, where each member $h\in{\cal H}$ is a function mapping ${\cal X}$ to ${\cal Y}$. Note that the hypothesis class ${\cal H}$ may have not be related to the model used for adversarial training in any way. Let $\ell:{\cal Y}\times{\cal Y}\rightarrow{\mathbb{R}}$ be a non-negative loss function. For every $h\in{\cal H}$, let function $f_{h}:{\cal X}\times{\cal Y}\rightarrow{\mathbb{R}}$ be defined by $f_{h}(x,y):=\ell(h(x),y)$. Let :=${\cal F}:=\{f_{h}:h\in{\cal H}\}$. From here on, we will restrict our attention to this “loss hypothesis class” ${\cal F}$ and study how well the members of ${\cal F}$ generalize from an i.i.d. training sample $\{(v_{i},y_{i})\}_{i=1}^{m}$ drawn from $\tilde{\mathcal{D}}_{\theta}$ to the unknown distribution $\tilde{\mathcal{D}}_{\theta}$. Specifically, for each $f\in{\cal F}$, the key quantity of our interest is the generalization gap $\left|\frac{1}{m}\sum\limits_{i=1}^{m}f(v_{i},y_{i})-\mathbb{E}_{(v,y)\sim\tilde{D}_{\theta}}f(v,y)\right|$.
@@
As we will soon show, the generalization gap turns out to be related to a key quantity characterizing a local property of the perturbation map $Q_{x,y,\theta}$, through which $\tilde{D}_{\theta}$ is induced.
Definition 1. Let ($\mathcal{X}^{\prime}$, $\|\cdot\|_{2}$) be a norm space equipped with the 2-norm. Given a map $T:\mathcal{X}\to\mathcal{X}^{\prime}$ and an arbitrary bounded measurable subset $C$ of ${\cal X}$, we define the $C$-dispersion of $T$ by
$\gamma_{C}(T):=\mathbb{E}_{x,x^{\prime}\sim\mathcal{U}(C)}\|T(x)-T(x^{\prime})\|_{2}^{2}$ (7)
where $\mathcal{U}(C)$ denotes a uniform distribution over $C$. Intuitively, this quantity measures on average how far two random points in $C$ spread after being mapped by $T$. Now restricting $T=Q_{x,y,\theta}$ and $C=\mathbb{B}(x,\epsilon)$, we have
$\gamma_{\mathbb{B}(x,\epsilon)}(\mathcal{Q}_{x,y,\theta})=\mathbb{E}_{\rho,\rho^{\prime}\sim\mathcal{U}([-\epsilon,+\epsilon]^{d})}\|\mathcal{Q}_{x,y,\theta}(x+\rho)-\mathcal{Q}_{x,y,\theta}(x+\rho^{\prime})\|_{2}^{2}.$

### 10 Feb 2024 v2

@@ Abstract
Adversarial training may be regarded as standard training with a modified loss function. But its generalization error appears much larger than standard training under standard loss. This phenomenon, known as robust overfitting, has attracted significant research attention and remains largely as a mystery. In this paper, we first show empirically that robust overfitting correlates with the increasing generalization difficulty of the perturbation-induced distributions along the trajectory of adversarial training (specifically PGD-based adversarial training). We then provide a novel upper bound for generalization error with respect to the perturbation-induced distributions, in which a notion of the perturbation operator, referred to “local dispersion”, plays an important role. Experimental results are presented to validate the usefulness of the bound and various additional insights are provided.
@@ 1 Introduction
Despite their outstanding performance, deep neural networks (DNNs) are known to be vulnerable to adversarial attacks where a carefully designed perturbation may cause the network to make a wrong prediction [1, 2]. Many methods have been proposed to improve the robustness of DNNs against adversarial perturbations [3, 4, 5], among which Projected Gradient Descend based Adversarial Training (PGD-AT) [3] is arguably the most effective [6, 7]. A recent work in [8] however revealed a surprising phenomenon in PGD-AT: after training, even though the robust error (i.e., error probability in the predicted label for adversarially perturbed instances) is nearly zero on the training set, it may remain very high on the testing set. For example, on the testing set of CIFAR10 [9], the robust error of PGD-AT trained model can be as large as 44.19%. This significantly contrasts the standard training: on CIFAR10, when the standard error (i.e., the error probability in the predicted label for non-perturbed instances) is nearly zero on the training set, its value on the testing set is only about 4%. This unexpected phenomenon arising in PGD-AT is often referred to as robust overfitting.
Figure 1: PGD-AT and the corresponding IDE results on the CIFAR-10 dataset with the Wide ResNet34 (WRN-34) model [10]. A significant increase in the IDE testing error is observed with the appearance of robust overfitting, suggesting a correlation between the generalization difficulty on $\tilde{\mathcal{D}}_{t}$ and robust overfitting.
@@ 3 Adversarial training and induced distributions
We consider a classification setting with input space ${\mathcal{X}}\subseteq\mathbb{R}^{d}$ and label space ${\mathcal{Y}}:=\{1,2,\cdots,K\}$. Let $\Theta$ be the parameter space of a model architecture of interest, and for each $\theta\in\Theta$, let $l_{\theta}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ denote a loss function (e.g, the cross-entropy loss) associated with the model $f_{\theta}$ with parameter $\theta$, where $\ell_{\theta}(v,y)=0$ if and only if $f_{\theta}(v)=y$.
For any data distribution $\mathcal{D}$ on $\mathcal{X}\times\mathcal{Y}$, the standard population error $R_{\cal D}(\theta)$ is defined as
$R_{\mathcal{D}}(\theta):=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[l_{\theta}(x,y)\right]$ (1)
and the adversarial population error $R_{\mathcal{D}}^{\rm adv}(\theta)$ is then
@@
where $\mathbb{B}(x,\epsilon)$ is customarily chosen as the $\infty$-norm ball centered at $x$ with radius $\epsilon$, or $\mathbb{B}(x,\epsilon):=\{t\in\mathbb{R}^{d}:\|t-x\|_{\infty}\leq\epsilon\}$. For later use, we will denote by ${\mathbb{U}}(x,\epsilon)$ the uniform distribution over ${\mathbb{B}}(x,\epsilon)$.
In adversarial training, the ultimate objective is to find a model parameter $\theta$ that minimizes $R_{\mathcal{D}}^{\rm adv}(\theta)$. Having no access to $\mathcal{D}$, in practice, a natural choice is to minimize the empirical version of $R_{\mathcal{D}}^{\rm adv}(\theta)$, namely,
$R_{S}^{\rm adv}(\theta):=\frac{1}{m}\sum\limits_{i=1}^{m}\max\limits_{v_{i}\in\mathbb{B}(x_{i},\epsilon)}l_{\theta}(v_{i},y_{i})$ (3)
on a training set $S:=\{(x_{i},y_{i})\}_{i=1}^{m}$ drawn i.i.d from $\mathcal{D}$. The most popular approach to solving this problem for neural networks is iterating between solving the inner maximization via $k$-step projected gradient descend (PGD) and updating $\theta$ through stochastic gradient descent. We now give a concise explanation of this procedure, referred to as PGD-AT, restating the procedure in [3].
@@
Here $\Pi_{\mathbb{B}(x,\epsilon)}:\mathbb{R}^{d}\to\mathbb{B}(x,\epsilon)$ denotes the operation of projecting onto the set $\mathbb{B}(x,\epsilon)$ and $\lambda\in\mathbb{R}_{+}$ is a hyperparameter. The sgn($\cdot$) function returns the sign of its input, and acts on a vector element-wise. The $k$-step PGD mapping ${\mathcal{Q}}_{x,y,\theta}:\mathbb{R}^{d}\to\mathbb{B}(x,\epsilon)$ is then
@@
We denote $\{{\cal Q}_{x,y,\theta}:(x,y)\in{\cal X}\times{\cal Y}\}$ collectively by ${\cal Q}_{\theta}$.
Iterations of PGD-AT At the $t^{\rm th}$ iteration of PGD-AT, where the model parameter is $\theta_{t}$, the solution of the inner maximization $\max_{v_{i}\in{\mathbb{B}}(x_{i},\epsilon)}l_{\theta_{t}}(v_{i},y_{i})$ is taken as $l_{\theta_{t}}({\cal Q}_{x_{i},y_{i},\theta_{t}}(x_{i}),y_{i})$, and the model parameter is updated, with learning rate $\eta$, by
$\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta_{t}}\left[\frac{1}{m}\sum\limits_{i=1}^{m}l_{\theta_{t}}\left({\cal Q}_{x_{i},y_{i},\theta_{t}}(x_{i}),y_{i}\right)\right]$ (6)
If $(X,Y)$ is drawn from ${\cal D}$, then we denote by $\tilde{\cal D}_{\theta_{t}}$, or simply $\tilde{\cal D}_{t}$, the distribution of $({\cal Q}_{X,Y,\theta_{t}}(X),Y)$, and refer to it as the induced distribution at iteration $t$. It is clear that the update equation in (6) is essentially one-step gradient descent of the standard notion of empirical risk on a training set drawn from $\tilde{\cal D}_{t}$. The dynamics of the gradient descent is however complicated by the fact that $\tilde{\cal D}_{t}$ continuously evolves with $t$, due to the dependency of $\mathcal{Q}_{\theta_{t}}$ on $\theta_{t}$. The thrust of this work is to investigate if the evolution trajectory of $\tilde{\cal D}_{t}$ along PGD-AT iterations may be related to robust overfitting.
@@ 4 Training on the induced distributions
The following experiments are conducted. First PGD-AT is performed on a training set $S$. Along this process, for a prescribed set of training iterations (or “checkpoints”) $I:=\{k_{1},k_{2},\ldots,k_{N}\}$, the model parameter $\theta_{t}$ at each checkpoint $t\in I$ is saved. Then at each checkpoint $t$, each example in the training set $S$ is perturbed according to ${\cal Q}_{\theta_{t}}$, giving rise to the perturbed training set $\tilde{S}_{t}$. The testing set $T$ is also similarly perturbed, giving rise to the perturbed testing set $\tilde{T}_{t}$. The model is then retrained fully on $\tilde{S}_{t}$, using standard training (i.e, without perturbation) with random initialization, where we denote the resulting learned model parameter by $\phi_{t}$. The model $\phi_{t}$ is then tested on $\tilde{T}_{t}$. Note that in this setting, both $\tilde{S}_{t}$ and $\tilde{T}_{t}$ are i.i.d. samples from $\tilde{\cal D}_{t}$. We call these experiments “induced distribution experiment” (IDE) for the ease of reference.
@@
Figure 2: The outcomes of additional experiments conducted on CIFAR-10. In the experiments, we perform PGD-AT with various weight decay rates and conduct IDEs for each of the PGD-AT variant. The blue curves are reproduced from Figure 1, serving as a reference for a clear comparison. The results further solidify the correlation between the robust overfitting and the IDE testing error.
@@
For each dataset, the experiments are repeated five times with different random seeds. The experimental results on CIFAR-10 and CIFAR-100 are shown in Figure 1, LABEL:fig:IDE_cifar100_&_resimg (a) and (b), where the green and yellow curves illustrate a phenomenon known as robust overfitting [8]: after a certain point in PGD-AT, the robust generalization gap (yellow curves), that is $|R_{\mathcal{D}}^{\rm adv}(\theta_{t})-R_{S}^{\rm adv}(\theta_{t})|$, steadily increases while the adversarial training error (green curves) constantly decreases. The red curves in the figures depict the testing errors of IDEs w.r.t various checkpoints in PGD-AT. Notably, a significant rise in the average IDE testing error is observed between the 80th and 120th checkpoints, increasing from 3.6% to 23.89% for CIFAR-10 and from 19% to 48.99% for CIFAR-100. Furthermore, this shift coincides with the onset of robust overfitting, where a significant rise in robust generalization gap is also observed. These results indicate that as $\theta_{t}$ evolves along PGD-AT, the model trained on the samples of $\tilde{\mathcal{D}}_{t}$ becomes harder to generalize. More importantly, this difficulty in generalization appears to be closely linked to the robust overfitting phenomenon. This potential connection is further demonstrated by the experimental results on the reduced ImageNet (see Figure LABEL:fig:IDE_cifar100_&_resimg (c)), where robust overfitting emerges at an earlier training stage and simultaneously a rise in the IDE testing error occurs. This increment in the IDE testing error is also substantial, with an averaged error of 21.65% at the 20th checkpoint elevating to 38.52% at the 60th checkpoint.
@@
Our experiments on MNIST (see Figure LABEL:fig:IDE_cifar100_&_resimg (d)) exhibits a scenario where a good robust generalization is achieved. Interestingly, a small IDE testing error is maintained throughout the evolution of $\tilde{\mathcal{D}}_{t}$ with the absence of robust overfitting. Figure 2 shows results from additional experiments on CIFAR-10. In these experiments, we perform PGD-AT with different level of weight decay to control the level of robust overfitting. Subsequently, IDEs are conducted for each such variant of PGD-AT. In Figure 2, each distinct color corresponds to a different weight decay factor utilized in PGD-AT. Within each color category, the dashed curves and the corresponding solid lines represent, respectively, the robust generalization gaps and the IDE results associated with that specific PGD-AT variant. From these results, we see that increasing the weight decay factor results in a notable reduction in the robust generalization gap, while conversely, decreasing the weight decay factor leads to the opposite effect. This is shown by the downward shift in the dashed curves across the three color categories. More noteworthy is a clear synchronization observed between each pair of dashed and solid curves, with lower dashed curves consistently corresponding to lower solid curves in the same color category.
All these results suggest a strong correlation between the IDE test error $R_{\tilde{\cal D}_{t}}(\phi_{t})$ and the robust generalization gap $R_{\cal D}^{\rm adv}(\theta_{t})-R_{S}^{\rm adv}(\theta_{t})$. Although by construction, $R^{\rm adv}_{\mathcal{D}}(\theta_{t})=R_{\tilde{\cal D}_{t}}(\theta_{t})$ and $R^{\rm adv}_{S}(\theta_{t})=R_{\tilde{S}_{t}}(\theta_{t})$, and hence the robust generalization gap can be re-expressed by $R_{\tilde{\cal D}_{t}}(\theta_{t})-R_{\tilde{S}_{t}}(\theta_{t})$, such a correlation is still quite surprising. This is because the learning of the parameter $\phi_{t}$ has been started from a completely random initialization and one would not expect the resulting parameter $\phi_{t}$ is linked to the parameter $\theta_{t}$ in any obvious way, despite that the latter contributes to shaping the distribution $\tilde{\cal D}_{t}$.
@@
A novel observation in this work, this correlation, is certainly curious in its own right and deserves further investigation. At this point, it has at least highlighted the impact of the dynamics of PGD-AT on robust overfitting, beyond the static quantities, such as loss landscape. It also suggests that the increasing difficulty of generalization inherent in the induced distribution plays an important role in robust overfitting.
@@ 5 Generalization properties of the induced distributions
There are two views of adversarial examples, “in-distribution view” (e.g., in [44, 45, 46]) and ”out-of-distribution view” (e.g., in [1, 47, 48]). In the out-of-distribution view, adversarial examples are considered as living outside the data manifold. In the in-distribution view, adversarial examples are considered as located within the support of the true data distribution but having low probability (density). This paper takes the in-distribution view and, instead of regarding the data distribution ${\cal D}$ as “absolutely robust”, we consider it only has a weaker notion of robustness guarantee. We now make precise this notion and our assumption of ${\cal D}$.
Underlying the data distribution ${\cal D}$, we assume that there exists another distribution ${\cal D}^{*}$ on ${\cal X}\times{\cal Y}$. We may re-express ${\cal D}^{*}$ as the pair $({\cal D}^{*}_{\cal X},h^{*})$, where ${\cal D}^{*}_{\cal X}$ is the marginal distribution of ${\cal D}^{*}$ on the input space ${\cal X}$ and $h^{*}$ is the ground-truth classifier, a function mapping ${\cal X}$ to ${\cal Y}$. The distribution ${\cal D}^{*}$ satisfies the following property: when $X$ is drawn from ${\cal D}^{*}_{\cal X}$ and $\rho$ is drawn independently of $X$ from ${\mathbb{U}}(0,\epsilon)$, $h^{*}(X+\rho)=h^{*}(X)$ with probability 1. That is, the classifier $h^{*}$ is absolutely robust (against adversarial perturbations with norm no larger than $\epsilon$) when $X$ is drawn from ${\cal D}^{*}_{\cal X}$—in this case, we also say, in short, “${\cal D}^{*}$ is absolutely robust”. We however note that in complex learning tasks, distributions with such strong notion of robustness may not arise naturally as the true data distribution. Instead, we assume that the true data distribution ${\cal D}$ is generated from ${\cal D}^{*}$ as follows: Draw $X$ from ${\cal D}^{*}_{\cal X}$, draw $\rho$ independent of $X$ from ${\mathbb{U}}(0,\epsilon)$, the distribution of $(X+\rho,h^{*}(X))$ is the true data distribution ${\cal D}$. That is, an instance drawn from the true data distribution is only guaranteed, with high probability, to resist all perturbations with magnitude no more than $\epsilon$. This assumption of data distribution ${\cal D}$ is consistent with the in-distribution perspective of adversarial examples and such weaker notion of robustness guarantee is also reminiscent of the notion of $(\epsilon,\delta)$-robustness recently presented in [49].
We now study the (standard) classification problem arising in the IDE setting at check point $t$, namely, that with $\tilde{\mathcal{D}}_{t}$ as the data distribution. Specifically, we wish to understand how difficult is such a learning task.
@@
Let ${\cal H}$ be a hypothesis class for this learning problem, where each member $h\in{\cal H}$ is a function mapping ${\cal X}$ to ${\cal Y}$. Note that the hypothesis class ${\cal H}$ may have not be related to the model used for adversarial training in any way. Let $\ell:{\cal Y}\times{\cal Y}\rightarrow{\mathbb{R}}$ be a non-negative loss function. For every $h\in{\cal H}$, let function $f_{h}:{\cal X}\times{\cal Y}\rightarrow{\mathbb{R}}$ be defined by $f_{h}(x,y):=\ell(h(x),y)$. Let :=${\cal F}:=\{f_{h}:h\in{\cal H}\}$. From here on, we will restrict our attention to this “loss hypothesis class” ${\cal F}$ and study how well the members of ${\cal F}$ generalize from an i.i.d. training sample $\{(v_{i},y_{i})\}_{i=1}^{m}$ drawn from $\tilde{\mathcal{D}}_{t}$ to the unknown distribution $\tilde{\mathcal{D}}_{t}$. Specifically, for each $f\in{\cal F}$, the key quantity of our interest is the generalization gap of each $f$ with respect to $\tilde{\cal D}_{t}$, defined by
${\rm GG}_{t}(f):=\left|\frac{1}{m}\sum\limits_{i=1}^{m}f(v_{i},y_{i})-\mathbb{E}_{(v,y)\sim\tilde{\cal D}_{t}}f(v,y)\right|.$
@@
As we will soon show, the generalization gap turns out to be related to a key quantity characterizing a local property of the perturbation map $Q_{x,y,\theta_{t}}$, through which $\tilde{\cal D}_{t}$ is induced.
Let $\rho$ and $\rho^{\prime}$ be drawn independently from ${\mathbb{U}}(0,\epsilon)$, and define
$\tilde{\gamma}_{t}(x,y):=\mathbb{E}_{\rho,\rho^{\prime}}\|\mathcal{Q}_{x,y,\theta_{t}}(x+\rho)-\mathcal{Q}_{x,y,\theta_{t}}(x+\rho^{\prime})\|_{2}^{2}.$ (7)